OpenStudy (anonymous):

Jean bought a bag of candy that contained red candies, green candies, and brown candies. Some candies are marked with a smile and others are not marked. The probabilities are displayed in the table below: Red Green Brown Smile 0.10 0.30 0.10 No smile 0.20 0.15 0.15 If a random candy is chosen from this bag, what is the probability that the candy was red or had no smile?

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

@Cosmichaotic could u help?

OpenStudy (anonymous):

Not sure on this one.. I'm horrible with statistics and probabilities

OpenStudy (anonymous):

@Brainybeauty @Kainui @anyoneelse plz help chris

OpenStudy (anonymous):

prob. of getting red candies = 0.10 prob. of getting candies with no smile = 0.20 + 0.15 + 0.15 = 0.5 Prob. of getting red candies + prob. of getting candies with no smile = 0.5 + 0.1 =0.6 Your answer is 0.6

OpenStudy (anonymous):

I imagined jellybeans while solving the answer though I like them more lol

OpenStudy (anonymous):

But don't worry the answer is right

OpenStudy (anonymous):

@chris!

OpenStudy (anonymous):

Ugh....thanks so much @Brainybeauty..... This problem still doesnt make sense to me. Like, why on earth does it say *or* instead of *and*?! It must be a glitch or typo lol? :P Anyway, just took the test, you were correct!!